OpenStudy (anonymous):

Solve each equation. Then show and check your work. 1.19=3y-5 2.8-5w=-37 3.0.4m-3=-1 4.n/3-8=-2 5.13=w/-3 -4

OpenStudy (anonymous):

1. 19+5=3y-5+5 24/3=3y/3 y=8